OpenStudy (anonymous):

Using complete sentences, explain the difference between an exponential function and a geometric series.(2 points)

OpenStudy (anonymous):

I have: They look same: exponential function: f(x) = 2^x geometric sequence: a[n] = 2^n the difference is x is a variable, while n is a positive number. So that exponential function's graph is continuous, geometric sequence is discrete.

OpenStudy (anonymous):

you don't explain the differences in the change in events.

OpenStudy (anonymous):

Can any body help me?